# Mean-field limits for Consensus-Based Optimization and Sampling

### v3

@@ 4.1 Well-posedness for the microscopic models
$\lvert\mathbf{X}\rvert:=\sqrt{\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}}.$
@@
$w_{j}(\mathbf{X}):=\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}.$
@@
where $\mathbf{X}_{t}:=\Bigl{(}X^{(1)}_{t},\dots,X^{(J)}_{t}\Bigr{)}$. The drift and diffusion coefficients are given by
$F(\mathbf{X}_{t}):=-\begin{pmatrix}X^{(1)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\\ \vdots\\ X^{(J)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\end{pmatrix}\in\mathbf{R}^{dJ},$
@@
$G(\mathbf{X}_{t}):=$ $\sqrt{2\lambda^{-1}}\operatorname{diag}\left(\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)},\dots,\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\right)\in\mathbf{R}^{dJ\times dJ}.$
@@
    $\forall\mathbf{X}\in\mathbf{R}^{dJ},\qquad\mathcal{L}\phi(\mathbf{X}):=F(\mathbf{X})\cdot\nabla\phi(\mathbf{X})+\frac{1}{2}\operatorname*{tr}\left(G^{\mathsf{T}}(\mathbf{X})\operatorname{D}^{2}\phi(\mathbf{X})G(X)\right)\leqslant c\phi(\mathbf{X}).$
Take $\phi(\mathbf{X}):=\tfrac{1}{2}|\mathbf{X}|^{2}$. Then
@@ 4.2 Well-posedness for the mean-field SDEs
We adapt the argument from [carrillo2018analytical], using the more general Assumption H and the improved stability estimates from Subsection 3.1. We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3. In the following, we endow the vector space $\mathcal{X}:=\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
$\left\lVert(m,\Gamma)\right\rVert:=\left({|m|^{2}+\left\lVert\Gamma\right\rVert^{2}_{\rm F}}\right)^{\frac{1}{2}}.$
@@
$C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}:=\left\{\varphi\colon[0,T]\rightarrow\mathcal{X}:\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}<\infty\right\},$
@@
$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}:=\sup_{t\in[0,T]}\bigl{\lVert}\varphi(t)\bigr{\rVert}+[\varphi]_{C^{0,\frac{1}{2}}}<\infty\qquad\text{and}\qquad[\varphi]_{C^{0,\frac{1}{2}}}:=\sup_{0\leqslant s<t\leqslant T}\frac{\bigl{\lVert}\varphi(t)-\varphi(s)\bigr{\rVert}}{|t-s|^{\frac{1}{2}}}.$
@@
admits a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$. Moreover, denoting by $\rho_{t}:=\operatorname{Law}(Y_{t})\in\mathcal{P}(\mathbf{R}^{d})$ the pointwise law of $Y_{t}$, the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belong to $C([0,T],\mathbf{R}^{d})$ and $C([0,T],\mathbf{R}^{d\times d})$ respectively.
@@
where $\rho_{t}:={\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3). Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined. Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5). The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [gilbarg1977elliptic, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
@@
$B_{R}:=\left\{(m,\Gamma)\in C([0,T],\mathcal{X}):\left\lVert(m,\Gamma)\right\rVert_{L^{\infty}([0,T])}\leqslant R\right\}.$
@@
It follows that $W_{p}(\rho_{t},\rho_{r})\leqslant L^{\frac{1}{p}}\lvert t-r\rvert^{\frac{1}{2}}$, which by Proposition 3.1 implies the Hölder continuity with exponent $\frac{1}{2}$ of the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$. This completes the proof that $\mathcal{T}$ is compact.
@@
$\mathrm{d}Y_{t}=-\Bigl{(}Y_{t}-\xi\mathcal{M}_{\beta}(\rho_{t})\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\xi\mathcal{C}_{\beta}(\rho_{t})}\,\mathrm{d}W_{t},\qquad\rho_{t}:={\rm Law}(Y_{t}).$
@@
For uniqueness, let $(m_{t},\Gamma_{t})$ and $(\widehat{m}_{t},\widehat{\Gamma}_{t})$ be two fixed points of $\mathcal{T}$ with corresponding solutions $Y_{t},\widehat{Y}_{t}\in C([0,T],\mathcal{X})$ of (4.3) and let $z_{t}:=Y_{t}-\widehat{Y}_{t}$. By definition of $\mathcal{T}$ and since $\mathcal{C}_{\beta}(\mu)$ is symmetric and positive semidefinite for all $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d})$, we have $\mathcal{R}\left(\Gamma_{t}\right)=\Gamma_{t}$ and $\mathcal{R}\bigl{(}\widehat{\Gamma}_{t}\bigr{)}=\widehat{\Gamma}_{t}$ for all $t\in[0,T]$. Let $\rho_{t},\widehat{\rho}_{t}\in\mathcal{P}(\mathbf{R}^{d})$ denote the marginal laws of $Y_{t}$ and $\widehat{Y}_{t}$, respectively. By the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$
@@
$\eta:=\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
@@
Using Proposition 3.1 and noting that $W_{p}(\rho_{r},\widehat{\rho}_{r})^{p}\leqslant\mathbf{E}\lvert z_{r}\rvert^{p}$ by definition of the Wasserstein distance, we conclude that
@@
where $\mathcal{L}_{u}\phi(y):=-\bigl{(}y-\mathcal{M}_{\beta}\left(\overline{\rho}_{u}\right)\bigr{)}\cdot\nabla\phi(y)+\lambda^{-1}\mathcal{C}_{\beta}\left(\overline{\rho}_{u}\right):\operatorname{D}^{2}\phi(y).$ It follows from (4.8) that $\overline{\rho}_{t}$ satisfies, for all smooth compactly supported $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$,
@@ 5.1 Proving mean-field results using stopping times
    $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sum_{j=1}^{J}\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
@@
    $\mathbf{E}\Bigl{[}W_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
@@ 5.2 Handling the case where $u>\ell>0$ in Assumption H
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
@@ A.1 Stability estimates for ratios of integrals
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert\leqslant CW_{p}(\mu,\nu).\qquad$
@@
Define $B_{K}:=\{x\in\mathbf{R}^{d}:\lvert x\rvert\leqslant K\}$ where $K:=2^{1/p}R$. By Markov’s inequality, we have
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\lvert\mathcal{S}(\mu)-\mathcal{S}(\nu)\rvert\leqslant L_{\mathcal{S}}W_{p}(\mu,\nu).$
@@
$\left(\iint\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{)}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant\sqrt{\frac{\mu[\lVert g(x)\rVert^{2}]}{\mu[h]}}\bigl{\lvert}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{\rvert}\leqslant\sqrt{\frac{C}{\eta}}L_{\mathcal{S}}W_{p}(\mu,\nu).$
@@
$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
@@ Conclusion
The claim follows then by combining the bounds on the three terms, and noting, for the first term, that $W_{p}(\mu,\nu)$ is bounded from above by the bracket on the right-hand side of (A.3) by definition of the Wasserstein distance. ∎
@@ A.2 Bound on weighted moments
for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved. The following is a generalization of [carrillo2018analytical, Lemma 3.3].
Proposition A.3.
Suppose that $f$ is bounded from below by $f_{\star}$ and that there are real numbers $u\geqslant\ell>0$ and positive constants $c_{u},c_{\ell}$ such that
$\forall x\in\mathbf{R}^{d},\qquad\left\{\begin{aligned} f(x)-f_{\star}\leqslant c_{u}\bigl{(}|x|^{u}+1\bigr{)}\\ f(x)-f_{\star}\geqslant c_{\ell}\bigl{(}|x|^{\ell}-1\bigr{)}.\end{aligned}\right.$ (A.6)
Then for all $p,q>0$ there are constants $C_{1}$ and $C_{2}$ depending on $(p,q,u,\ell,c_{u},c_{\ell})$ such that
The proof relies on the following lemma.
@@
$\lambda:=\left(\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{1}{q}}<\infty.$
@@
$F(\gamma):=\frac{\int_{\mathbf{R}^{+}}g(z)\,\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}$
@@
$F^{\prime}(\gamma)=-\int_{\mathbf{R}^{+}}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z),\qquad L_{\gamma}\pi(\mathrm{d}z):=\frac{\operatorname{e}^{-\gamma z}\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)},\qquad\overline{z}:=\int_{\mathbf{R}^{+}}z\,L_{\gamma}\pi(\mathrm{d}z).$
@@
Proof of Proposition A.3.
@@
$\lvert x\rvert^{p}=\left(\lvert x\rvert^{\ell}\right)^{\frac{p}{\ell}}\leqslant\left(\frac{f(x)}{c_{\ell}}+1\right)^{\frac{p}{\ell}}.$ (A.10)
For convenience, we denote by $C$ any constant depending only on $(p,q,\ell,u,c_{\ell},c_{u})$. From (A.10) we deduce that

### v4

@@ 4.1 Well-posedness for the microscopic models
$\lvert\mathbf{X}\rvert\coloneq\sqrt{\sum_{j=1}^{J}\left\lvert X^{(j)}\right\rvert^{2}}.$
@@
$w_{j}(\mathbf{X})\coloneq\frac{\operatorname{e}^{-\beta f\left(X^{(j)}\right)}}{\sum_{k=1}^{J}\operatorname{e}^{-\beta f\left(X^{(k)}\right)}}.$
@@
where $\mathbf{X}_{t}\coloneq\Bigl{(}X^{(1)}_{t},\dots,X^{(J)}_{t}\Bigr{)}$. The drift and diffusion coefficients are given by
$F(\mathbf{X}_{t})\coloneq-\begin{pmatrix}X^{(1)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\\ \vdots\\ X^{(J)}_{t}-\mathcal{M}_{\beta}\left(\mu^{J}_{t}\right)\end{pmatrix}\in\mathbf{R}^{dJ},$
@@
$G(\mathbf{X}_{t})\coloneq$ :- $\sqrt{2\lambda^{-1}}\operatorname{diag}\left(\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)},\dots,\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{t}\right)}\right)\in\mathbf{R}^{dJ\times dJ}.$
@@
    $\forall\mathbf{X}\in\mathbf{R}^{dJ},\qquad\mathcal{L}\phi(\mathbf{X})\coloneq F(\mathbf{X})\cdot\nabla\phi(\mathbf{X})+\frac{1}{2}\operatorname*{tr}\left(G^{\mathsf{T}}(\mathbf{X})\operatorname{D}^{2}\phi(\mathbf{X})G(X)\right)\leqslant c\phi(\mathbf{X}).$
Take $\phi(\mathbf{X})\coloneq\tfrac{1}{2}|\mathbf{X}|^{2}$. Then
@@ 4.2 Well-posedness for the mean-field SDEs
We adapt the argument from [carrillo2018analytical], using the more general Assumption H and the improved stability estimates from Subsection 3.1. We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3. In the following, we endow the vector space $\mathcal{X}\coloneq\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
$\left\lVert(m,\Gamma)\right\rVert\coloneq\left({|m|^{2}+\left\lVert\Gamma\right\rVert^{2}_{\rm F}}\right)^{\frac{1}{2}}.$
@@
$C^{0,\frac{1}{2}}\bigl{(}[0,T],\mathcal{X}\bigr{)}\coloneq\left\{\varphi\colon[0,T]\rightarrow\mathcal{X}:\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}<\infty\right\},$
@@
$\left\lVert\varphi\right\rVert_{C^{0,\frac{1}{2}}}\coloneq\sup_{t\in[0,T]}\bigl{\lVert}\varphi(t)\bigr{\rVert}+[\varphi]_{C^{0,\frac{1}{2}}}<\infty\qquad\text{and}\qquad[\varphi]_{C^{0,\frac{1}{2}}}\coloneq\sup_{0\leqslant s<t\leqslant T}\frac{\bigl{\lVert}\varphi(t)-\varphi(s)\bigr{\rVert}}{|t-s|^{\frac{1}{2}}}.$
@@
admits a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$. Moreover, denoting by $\rho_{t}\coloneq\operatorname{Law}(Y_{t})\in\mathcal{P}(\mathbf{R}^{d})$ the pointwise law of $Y_{t}$, the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ belong to $C([0,T],\mathbf{R}^{d})$ and $C([0,T],\mathbf{R}^{d\times d})$ respectively.
@@
where $\rho_{t}\coloneq{\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3). Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined. Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5). The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [gilbarg1977elliptic, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
@@
$B_{R}\coloneq\left\{(m,\Gamma)\in C([0,T],\mathcal{X}):\left\lVert(m,\Gamma)\right\rVert_{L^{\infty}([0,T])}\leqslant R\right\}.$
@@
It follows that $\mathcal{W}_{p}(\rho_{t},\rho_{r})\leqslant L^{\frac{1}{p}}\lvert t-r\rvert^{\frac{1}{2}}$, which by Proposition 3.1 implies the Hölder continuity with exponent $\frac{1}{2}$ of the functions $t\mapsto\mathcal{M}_{\beta}(\rho_{t})$ and $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$. This completes the proof that $\mathcal{T}$ is compact.
@@
$\mathrm{d}Y_{t}=-\Bigl{(}Y_{t}-\xi\mathcal{M}_{\beta}(\rho_{t})\Bigr{)}\,\mathrm{d}t+\sqrt{2\lambda^{-1}\xi\mathcal{C}_{\beta}(\rho_{t})}\,\mathrm{d}W_{t},\qquad\rho_{t}\coloneq{\rm Law}(Y_{t}).$
@@
For uniqueness, let $(m_{t},\Gamma_{t})$ and $(\widehat{m}_{t},\widehat{\Gamma}_{t})$ be two fixed points of $\mathcal{T}$ with corresponding solutions $Y_{t},\widehat{Y}_{t}\in C([0,T],\mathcal{X})$ of (4.3) and let $z_{t}\coloneq Y_{t}-\widehat{Y}_{t}$. By definition of $\mathcal{T}$ and since $\mathcal{C}_{\beta}(\mu)$ is symmetric and positive semidefinite for all $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d})$, we have $\mathcal{R}\left(\Gamma_{t}\right)=\Gamma_{t}$ and $\mathcal{R}\bigl{(}\widehat{\Gamma}_{t}\bigr{)}=\widehat{\Gamma}_{t}$ for all $t\in[0,T]$. Let $\rho_{t},\widehat{\rho}_{t}\in\mathcal{P}(\mathbf{R}^{d})$ denote the marginal laws of $Y_{t}$ and $\widehat{Y}_{t}$, respectively. By the Burkholder–Davis–Gundy inequality, we have for all $t\in[0,T]$
@@
$\eta\coloneq\inf\left\{\lambda_{\rm min}\Bigl{(}\mathcal{C}_{\beta}(\rho_{t})\Bigr{)}:t\in[0,T]\right\}>0.$
@@
Using Proposition 3.1 and noting that $\mathcal{W}_{p}(\rho_{r},\widehat{\rho}_{r})^{p}\leqslant\mathbf{E}\lvert z_{r}\rvert^{p}$ by definition of the Wasserstein distance, we conclude that
@@
where $\mathcal{L}_{u}\phi(y)\coloneq-\bigl{(}y-\mathcal{M}_{\beta}\left(\overline{\rho}_{u}\right)\bigr{)}\cdot\nabla\phi(y)+\lambda^{-1}\mathcal{C}_{\beta}\left(\overline{\rho}_{u}\right):\operatorname{D}^{2}\phi(y).$ It follows from (4.8) that $\overline{\rho}_{t}$ satisfies, for all smooth compactly supported $\phi\colon\mathbf{R}^{d}\to\mathbf{R}$,
@@ 5.1 Proving mean-field results using stopping times
    $\Omega_{J}:=\left\{\omega\in\Omega:\frac{1}{J}\sup_{t\in[0,T]}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}(\omega)\right\rvert^{p}\geqslant R\right\}.$ (5.1)
@@
    $\mathbf{E}\Bigl{[}\mathcal{W}_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
@@ 5.2 Handling the case where $u>\ell>0$ in Assumption H
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant C_{\varepsilon}J^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
@@ A.1 Stability estimates for ratios of integrals
$\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\left\lVert\frac{\mu[g]}{\mu[h]}-\frac{\nu[g]}{\nu[h]}\right\rVert\leqslant C\mathcal{W}_{p}(\mu,\nu).\qquad$
@@
Define $B_{K}\coloneq\{x\in\mathbf{R}^{d}:\lvert x\rvert\leqslant K\}$ where $K\coloneq 2^{1/p}R$. By Markov’s inequality, we have
@@
    $\forall(\mu,\nu)\in\mathcal{P}_{p,R}(^{d})\times\mathcal{P}_{p,R}(^{d}),\qquad\lvert\mathcal{S}(\mu)-\mathcal{S}(\nu)\rvert\leqslant L_{\mathcal{S}}\mathcal{W}_{p}(\mu,\nu).$
@@
$\left(\iint\left\lVert\frac{g(x)}{\sqrt{\mu[h]}}\bigl{(}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{)}\right\rVert^{2}\,\pi(\mathrm{d}x\,\mathrm{d}y)\right)^{1/2}\leqslant\sqrt{\frac{\mu[\lVert g(x)\rVert^{2}]}{\mu[h]}}\bigl{\lvert}\mathcal{S}(\mu)-\mathcal{S}(\nu)\bigr{\rvert}\leqslant\sqrt{\frac{C}{\eta}}L_{\mathcal{S}}\mathcal{W}_{p}(\mu,\nu).$
@@
$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}\mathcal{W}_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
@@ Conclusion
The claim follows then by combining the bounds on the three terms, and noting, for the first term, that $\mathcal{W}_{p}(\mu,\nu)$ is bounded from above by the bracket on the right-hand side of (A.3) by definition of the Wasserstein distance. ∎
@@ A.2 Bound on weighted moments
for a constant $C$ depending only on $\beta$ and $f$. When $f$ is unbounded, proving an inequality of this type is more involved.
To this end, we first state and prove the following auxiliary lemma.
@@
$\lambda\coloneq\left(\int_{\mathbf{R}^{+}}y^{q}\,\nu(\mathrm{d}y)\right)^{\frac{1}{q}}<\infty.$
@@
$F(\gamma)\coloneq\frac{\int_{\mathbf{R}^{+}}g(z)\,\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)}$
@@
$F^{\prime}(\gamma)=-\int_{\mathbf{R}^{+}}g(z)\left(z-\overline{z}\right)\,L_{\gamma}\pi(\mathrm{d}z),\qquad L_{\gamma}\pi(\mathrm{d}z)\coloneq\frac{\operatorname{e}^{-\gamma z}\pi(\mathrm{d}z)}{\int_{\mathbf{R}^{+}}\operatorname{e}^{-\gamma z}\,\pi(\mathrm{d}z)},\qquad\overline{z}\coloneq\int_{\mathbf{R}^{+}}z\,L_{\gamma}\pi(\mathrm{d}z).$
@@
Armed with this result, we are ready to prove the following generalization of [carrillo2018analytical, Lemma 3.3].
Proposition A.4.
Suppose that $f$ is bounded from below by $f_{\star}$ and that there are real numbers $u\geqslant\ell>0$ and positive constants $c_{u},C_{u},c_{\ell},C_{\ell}$ such that
$\forall x\in\mathbf{R}^{d},\qquad\left\{\begin{aligned} f(x)-f_{\star}\leqslant c_{u}|x|^{u}+C_{u}\\ f(x)-f_{\star}\geqslant c_{\ell}|x|^{\ell}-C_{\ell}.\end{aligned}\right.$ (A.8)
Then for all $p,q>0$ there are constants $C_{1}$ and $C_{2}$ depending on $(p,q,u,\ell,c_{u},C_{u},c_{\ell},C_{\ell})$ such that
Proof.
@@
$\lvert x\rvert^{p}=\left(\lvert x\rvert^{\ell}\right)^{\frac{p}{\ell}}\leqslant\left(\frac{f(x)+C_{\ell}}{c_{\ell}}\right)^{\frac{p}{\ell}}.$ (A.10)
For convenience, we denote by $C$ any constant depending only on $(p,q,\ell,u,c_{\ell},C_{\ell},c_{u},C_{u})$. From (A.10) we deduce that